# Learning Coherent Matrixized Representation in Latent Space for Volumetric 4D Generation

### v2

@@ 3.3 Spatio-Temporal Diffusion for Conditional 4D Volumetric Generation
$\mathcal{L}_{LDM}=\mathop{\mathbb{missing}}{E}_{\mathcal{M},f,\epsilon,t}\left[{\epsilon-\epsilon_{\theta}(\mathcal{M}_{t},t,F\{D\circ\tau_{i}(f_{i})\})}^{2}\right],$ (14)
where $D$ is a dropout operation enabling classifier-free guidance, $\tau_{i}(f_{i})$ is the $i$-th modality feature encoded by the task-specific encoder $\tau_{i}$, and $F$ refers to a simple concatenation. This approach not only ensures uniformity in the denoising process but also significantly reduces computational overhead.
@@ 4 Experiments
Data Preparation: For 3D geometry evaluation, we use car, chair, and airplane categories from ShapeNet [7] (details in supplementary). For colored 3D objects evaluation, we use T-pose objects from 3DBiCar [28]. We use 16-frame animal sequences from DeformingThings4D (DT4D) [23] ($1772$ sequences) and 20-frame colored cartoon sequences from 3DBiCar (sample $30$k sequences via latent space interpolation) for both unconditional and conditional generation. We also use $110$k subset and $323$k of Objaverse [11] for conditional 4D generation.
Evaluation Metrics: We use MMD, COV, and 1-NNA for geometry evaluation, following [27, 52]. Lower MMD, higher COV, and 50% 1-NNA are optimal. For 4D sequence evaluation, we average 3D metrics across frames. To assess inter-frame consistency, we compute CD-VAR and COV-VAR by calculating Chamfer Distance and COV variance across frames. For color evaluation, PSNR and SSIM are computed by sampling 16,384 surface points multiple times. We also use CLIP score (CLIP-S).
Implementation Details: The loss weights $\lambda_{1}$ and $\lambda_{2}$ are set to $1$. More details about dataset, settings, results and analysis are in supplementary.
Category $U_{\phi}$ MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
$\times$ 0.012 0.007 69 2.791 36
3DBiCar $\checkmark$ 0.009 0.004 73 0.834 41
$\times$ 0.037 0.003 69 3.179 36
DT4D $\checkmark$ 0.030 0.001 76 2.91 43
Table 1: Ablation study on 4D sequence generation: $U_{\phi}$ enhances SDF details, improving shape evaluation.
@@ 4.1 Ablation Studies
Refinement Module $U_{\phi}$. The refinement U-Net $U_{\phi}$ improves the quality of coarse SDF with a 3D U-Net. Tab. 1 shows ablation study of $U_{\phi}$ in 4D sequence generation, where $U_{\phi}$ improves all metrics because $U_{\phi}$ enhances the geometric performance of all frames within the 4D sequence, thereby indirectly promoting consistency across frames. Note that although $U_{\phi}$ improves the quality of the learned SDF, the inherent limitations of the SDF representation suggest that utilizing a more advanced 3D representation will further enhance the performance of individual objects.
HCSTA module. Tab. 2 shows ablation studies on attention modulation, order, and unification. We first evaluate attention for shape-color (S), conditioned (C), and temporal (T) dimensions. Tab. 2 (rows 1-3) shows shape-color cross-attention contributes most. In rows 4-7, we permute the S, C, T orders, the S-C-T order performing best. Finally, replacing the proposed hierarchical attention with a unified strategy (shared attention) leads to a significant drop (rows 8-9), our method outperforms it.
Type S C T MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
Module $\times$ $\checkmark$ $\checkmark$ 0.014 0.018 61 2.877 33
$\checkmark$ $\times$ $\checkmark$ 0.010 0.015 68 2.165 39
$\checkmark$ $\checkmark$ $\times$ 0.011 0.011 71 2.663 38
Order S-C-T 0.009 0.007 73 1.834 41
S-T-C 0.010 0.011 71 2.165 39
C-T-S 0.012 0.015 68 2.663 38
C-S-T 0.011 0.014 69 2.731 35
Uni-Stra. $\checkmark$ 0.013 0.018 67 3.002 34
$\times$ 0.009 0.007 73 1.834 41
Table 2: Ablation on attentions modulation, orders, and unifying.
L1 R SSIM F PSNR SSIM
$\checkmark$ $\checkmark$ $\times$ $\times$ 29.15 0.79
$\checkmark$ $\times$ $\checkmark$ $\times$ 26.37 0.72
$\times$ $\checkmark$ $\checkmark$ $\times$ 28.76 0.75
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 9.64 0.64
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 34.51 0.85
Table 3: Quantitative results of color ablation study.
Coherent 3D Shape and Color Modeling. We evaluate the impact of L1 loss, remeshing (R), structural similarity loss (SSIM), and a shape-agnostic approach (F) that assigns colors from a uniform distribution per object (see Suppl. for details). As evidenced in Tab. 3, integrating L1, R, and SSIM optimizes color fidelity: L1 ensures foundational accuracy, SSIM improves local structural coherence, and R enhances vertex density for refined color detail. Our coherent 3D shape and color modeling constructs a continuous color field adaptive to varied geometries, ensuring shape-constrained color continuity and boosting accuracy. The combined R and SSIM metrics further enable precise capture and learning of color details.
Category Att. MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
$\times$ 0.011 0.018 69 2.988 36
3DBiCar $\checkmark$ 0.009 0.007 73 1.834 41
$\times$ 0.034 0.015 73 3.77 36
DT4D $\checkmark$ 0.030 0.001 76 2.91 43
Table 4: Ablation study of HCSTA for 4D generation.
Inter-frame Consistency. We conduct a comparison between the HCSTA layer and the standard attention mechanism utilized in traditional latent diffusion models. As shown in Tab. 4, the HCSTA layer significantly enhances temporal consistency, as reflected by improvements across all metrics. This enhancement is due to the HCSTA layer’s ability to effectively capture the frame-to-frame evolution of $\mathcal{M}$, resulting in smoother transitions between frames.
Decoupled Color and Shape. We validate the efficacy of our decoupled shape and color framework by assessing fitting quality for combined and decoupled latents in Tab. 5. Conditional 4D generation quality and shape-color consistency are further evaluated through a user study (Tabs. 56), where participants compared our method to baselines across three metrics, selecting the superior approach per vote. Fig. 5 demonstrates that the decoupled representation allows color editing without altering shape. Results consistently show that ours improves generation quality and enables independent color modifications, with decoupled representations preventing shape information disruption during color reconstruction, enhancing fitting performance.
Figure 5: Visualization of color editing without shape alteration.
Method MMD COV(%) PSNR SSIM CLIP-S
Combined 0.021 57 29.13 0.72 0.801
Decoupled 0.009 73 34.51 0.85 0.911
Table 5: Quantitative comparison of fitting quality between combined and decoupled shape and color representations.
Metric Combined Decoupled
Color and Shape Consistency 7.33% 92.67%
Interframe Consistency 6.31% 93.69%
Table 6: User study voting rates for consistency of color and shape, and interframe consistency.
Method MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
Voxel 0.293 0.035 30 1.857 91
Hyper. 0.044 0.011 52 2.138 69.3
Ours 0.009 0.007 73 1.834 41
Table 7: 4D sequence generation comparison on 3DBiCar dataset.
4.2 Unconditional 4D Sequence Shape Generation
We demonstrate our model’s capability to generate diverse objects and motions in unconditional 4D generation. Additionally, we showcase its ability to produce coherent sequences through latent matrixized 4D representation interpolation. Our evaluation, which focuses on shape and temporal consistency, is conducted using the 3DBiCar and DeformingThings4D datasets. Tab. 7 presents quantitative comparisons of 4D sequence generation, demonstrating that our model consistently outperforms the voxel baseline and HyperDiffusion [13] (Hyper.) across all evaluated metrics. These alternative methods face challenges related to voxel quality and high-dimensional sparsity.
Method Speed Color CLIP-S
3DBiCar Obj.-110k Obj.-323k
DG4D [35] 6.5 min $\checkmark$ 0.803 0.790 0.785
Hyper. [13] 0.54s $\times$ - - -
STAG4D [48] 7.5 min $\checkmark$ 0.908 0.895 0.906
L4GM [36] 3.7s $\checkmark$ 0.940 0.925 0.930
4Diffusion [49] 46.3s $\checkmark$ 0.909 0.911 0.903
Ours 0.625s $\checkmark$ 0.911 0.905 0.915
Table 8: Comparison of 4D generation method in speed, and quality on 3DBiCar, Objaverse-110k and Objaverse-323k.
Fig. 6 explores motion diversity and interpolation, with each row depicting a different action performed by the same object. This demonstrates our model’s capability to generate varied motions for the same object with high quality, highlighting both its versatility and its detail-rich output. Additionally, the 2nd, 4th, and 6th columns display objects decoded from interpolated $\mathcal{M}$. A comparison of these columns with their adjacent ones reveals that the interpolated results exhibit strong temporal consistency, indicating that our latent matrixized 4D representation supports the generation of sequences with variable-length frame density through interpolation.
4.3 Conditonal 4D Sequence Generation
Figure 9: Generated volumetric 4D sequence.
Figure 10: Identity transfer experiments.
Our model uses diffusion to generate latent 4D sequence representation $\mathcal{M}$ from image and text conditions. First, we compare it with DreamGaussian4D (DG4D) [35], HyperDiffusion [13], STAG4D [48], L4GM [36], and 4Diffusion [49] in Tab. 8 using 3DBiCar and Objaverse 110k and 323k subsets. DG4D supports colored 4D sequences but is slow and memory-intensive. Although STAG4D significantly enhances output quality, this improvement is accompanied by increased computational overhead. HyperDiffusion, while fast, compromises on both color and geometric quality. Although L4GM and 4Diffusion achieve similar generation quality to ours, they require significantly more time and memory. In contrast, our method balances speed and memory efficiency in generation. We also test the spatiotemporal consistency and the alignment with the input. Please refer to appendix for more details.
Generation via conditions. First, we present our conditional 4D sequence generation results in Fig. 7. We include images from different viewing angles and depth maps of the generated sequences to demonstrate our model’s ability to produce plausible, topologically continuous, and visually consistent 4D sequences. Fig. 7 illustrates that our model effectively learns color and shape information, and generates 4D motion sequences from image and text prompts. By balancing the trade-off between 3D shape quality and sequence coherence, our method generates the 4D representation $\mathcal{M}$, which can be accurately decoded into 4D sequences maintaining diversity and alignment with image-text guidances.
Geometric quality comparison. We compare our method with DG4D in Fig. 11, providing color images and normal maps from different views to highlight the differences. Due to GPU limitations, DG4D exhibits poor rendering quality outside supervised views. In contrast, our model generates 4D sequences with more complete appearances and smoother normal and depth maps. Our approach outperforms DG4D by balancing memory efficiency with high-quality sequence generation, focusing on accurate shape representation for each identity, while DG4D emphasizes rendering effects of Gaussian primitives without a precise 3D representation.
Variable-length 4D sequences. To demonstrate the capability of generating arbitrarily long 4D sequences, we create latent 4D representations of the same object performing different actions using varying motion prompts, then apply interpolation to generate intermediate frames. Fig. 8 shows that our method can produce variable-length 4D sequences through identity transfer and interpolation. This validates the practicality of our approach for generating coherent 4D sequences in stages and interpolations.
Volumetric property. To demonstrate the volumetric property of our generated results, we create volumetric 4D sequences, place them in a shared scene, and perform free navigation and rendering. Results in Fig. 9 show that our method supports navigation across different views (a), (b), and (c) while maintaining high-quality rendering. Additionally, our approach enables free editing and manipulation of each frame, distinguishing it from other free-form 4D generation models.
Figure 11: Qualitative comparison with DreamGaussian4D.
Cross Identity transfer. To evaluate the controllability of our method, we apply the same motion prompt to images of different identities, generating 4D sequences with consistent motion but varying identities. Results in Fig. 10 show that our method retains the input’s identity and appearance while accurately capturing the motion described in the prompt. This demonstrates that our latent matrixized 4D sequence representation effectively decouples shape, color, and motion, with the HCSTA layer controlling the diffusion generation process.
5 Conclusion
We propose an image-text conditioned volumetric 4D sequence generation method balancing 3D shape quality with sequence coherence, while maintaining diversity and alignment with input guidances. Experiments demonstrate our method’s efficiency and superior performance. We hope it will be a useful attempt that opens up new possibilities for the representation and generation of complex 4D sequences. Limitations. The lack of large-scale real-world datasets limits 4D generation performance. Improved 3D/4D representations and the use of generative priors will be key areas for future breakthroughs.
[…]
Supplementary Material
Figure 12: Overview of generated volumetric 4D sequence.
6 Diffusion Process
Due to our decoupled design of shape and color in our latent 4D representation, we can denote the latent representation as =$\mathcal{M}\in=\{\{\mathbf{z}_{i}\in\mathbb{R}^{d_{1}}\},\{\mathbf{c}_{i}\in\mathbb{R}^{d_{2}}\}\}$, consisting $N$ frames with shape latent vectors $\{\mathbf{z}_{i}\in\mathbb{R}^{d_{1}}\}$ and color latent vectors $\{\mathbf{c}_{i}\in\mathbb{R}^{d_{2}}\}$. In our HSCTA layer, the diffusion process can be described as follows: The shape-color cross-attention layer enables the color latent codes to be supervised by the shape latent codes:
$\min_{\theta}\mathbb{E}[\sum_{t=1}^{T}\log p_{\theta}(\mathbf{c}_{t-1}|\mathbf{c}_{t},\mathbf{z}_{t})],$ (15)
Then the condition are injected to control the generation of $\mathbf{c}_{t-1}$ and $\mathbf{z}_{t-1}$:
$\min_{\theta}\mathbb{E}[\sum_{t=1}^{T}\log p_{\theta_{2}}(\mathbf{c}_{t-1},\mathbf{z}_{t-1}|\mathbf{c}_{t},\mathbf{z}_{t},\{f_{i}^{T}\circ f_{i}^{I}\})],$ (16)
Finally, we perform the temporal self-attention layer, to enhance spatio-temporal consistency:
$\begin{gathered}\min_{\theta}\mathbb{E}[\sum_{t=1}^{T}\log p_{\theta}(\mathbf{z}_{t-1}^{i+1},\mathbf{c}_{t-1}^{i+1}|\mathbf{z}_{t-1}^{i},\mathbf{c}_{t-1}^{i})]\end{gathered}$ (17)
7 Experiments Details
For the shape auto-decoder $D_{\theta}$, we employ an 8-layer MLP to decode the SDF values, using a resolution of $64\times 64\times 64$ to construct a 6-layer 3D U-Net $U_{\phi}$ on DT4D and $128\times 128\times 128$ on 3DBiCar and ShapeNet. For the color auto-decoder $C_{\theta}$, we use an 8-layer MLP. For denoising network, we use an 8-layer UNet augmented with HCSTA Layer.
We provide an overview of the hyperparameters of our models in Tab. 9, including the size of the latent 4D representation $\mathcal{M}$ and loss weights. We evaluated the metrics and ultimately selecting the combination $\{\mathcal{M}\in\mathbb{R}^{N\times(256+256)},~{}\lambda_{1}=1.0,~{}\lambda_{2}=1.0\}$. Balancing the L1 and SSIM losses aids the model in learning both the primary colors and the finer color details in a more balanced manner. Moreover, experimentation with dimensions larger than the current setting (256+256) was conducted. However, it was observed that when the scale of $\mathcal{M}$ exceeds 256 (as indicated in row 3 of Tab. 9), the data distribution becomes too sparse to support the diffusion generation for volumetric 4D sequence effectively. Finally, analysis of the remaining parameter settings (rows 4-5) reveals that an unbalanced weighting of loss terms significantly leads to performance degradation.
@@ 7 Experiments Details
Table 9: Hyperparameters for the conditional models trained on 3DBiCar.
We follow prior works in evaluating Minimum Matching Distance (MMD), Coverage (COV), and 1-Nearest-Neighbor Accuracy (1-NNA). For MMD, lower is better; for COV, higher is better; for 1-NNA, 50% is the optimal.
$\text{COV}(S_{g},S_{r})=\frac{|\{\arg\min_{Y\in S_{r}}D(X,Y)|X\in S_{g}\}|}{|S_{r}|}$ (19)
@@
where in the 1-NNA metric $N_{X}$ is a point cloud that is closest to $X$ in both generated and reference dataset, i.e.,
$N_{X}=\arg\min{K\in S_{r}\cup S_{g}}D(X,K)$
We use a Chamfer Distance (CD) distance measure $D(X,Y)$ for computing these metrics in 3D, and report CD values multiplied by a constant $10^{2}$. To evaluate 4D shapes, we extend to the temporal dimension for $T$ frames:
$D(X,Y)=\frac{1}{T}\sum_{t=0}^{T-1}CD(X[t],Y[t]).$
All videos are rendered from Blender (see Fig. 12), mesh sequences are imported by Stop-Motion-Obj. Stop-Motion-OBJ is a tool that lets users import a volumetric 4D sequence, play them back in real time, then render them out to an animation. Each mesh may have a different vertex count or poly count. This is especially useful for importing fluid simulations, for visualization of scientific data, and for rendering of volumetric 4D scan data. Benifit from the volumetric 4D sequence, our method also supports switching between rendering with color, white model rendering, and normal rendering, while providing high-quality results (see Fig. 13).
Figure 13: Render results in normal and w/o color mode.
8 Additional Evaluation
First, we show the visualization results for ablation study on $U_{\phi}$ in Fig. 14. The refinement module $U_{\phi}$ learns more geometry details compared to a single decoder $D_{\theta}$. Then We compare 3D shape reconstruction quality among DeepSDF [31], SDFusion [9], and our method. SDFusion uses a 3D latent diffusion model with a VQVAE decoder for shape generation and a 2D diffusion model for texture rendering. DeepSDF decodes SDF values without color using an MLP decoder. We evaluate geometry accuracy on ShapeNet (car, chair, airplane) and 3DBiCar (T-pose) datasets. Tab. 10 shows our method outperforms SDFusion and DeepSDF in geometry generation and surpasses SDFusion in color generation, confirming our coherent 3D shape and color modeling learns more shape details.
@@ 8 Additional Evaluation
Table 10: Quantitative comparison of Single Shape. Our method surpasses other models across the majority of evaluation metrics.
Figure 14: Illustration of 3D shape ablation study.
We provide the full results of evaluating our model’s ability to generate 4D object sequences on the 3DBiCar and DeformingThings4D datasets. Our model consistently delivers superior results across all metrics (Tab. 11)
Category Method MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
Voxel 0.293 0.035 30 1.857 91
Hyper. 0.044 0.011 52 2.138 69.3
3DBiCar Ours 0.009 0.007 73 1.834 41
Voxel 0.219 0.044 35 1.673 85
Hyper. 0.155 0.013 45 1.735 62
DT4D Ours 0.030 0.001 76 2.91 43
Table 11: Quantitative comparison of 4D Sequence Generation.
Figure 15: Conditions for 4D generation.
9 Data Preparation
DeformingThings4D [23] is a synthetic dataset containing 1,972 animation sequences spanning 31 categories of humanoids and animals. However, the lack of color information and image annotations in DeformingThings4D makes it challenging to use for conditional 4D sequence generation.
The 3DBiCar [28] dataset provides a script for sampling high-quality, topologically consistent 3D textured models of biped cartoon characters, manually crafted by professional artists, covering diverse identities, shapes, and textures across $15$ species and $4$ image styles. Sourced from Internet images, it supports applications such as single-view reconstruction, sketch-based modeling, and 3D animation. While providing mask and origin images for 4D sequence design, it lacks text captions. We employed Large Language Models (LLMs) for initial image descriptions, manually refined for accuracy in shape, color, and pose, with plans to release these captions post-acceptance. For conditioning, masked images were derived from provided masks and replicated across frames, while text conditions incorporated frame-order prefixes (e.g., ”$n$-th frame”) and were encoded with CLIP alongside image conditions (examples in Fig. 15). Training data included T-pose to target-pose transitions for model optimization.
Figure 16: Pipeline of non-unified color learning. Unlike the proposed representation, we invert the marching cubes algorithm to locate reconstructed vertices tied to query points, termed key query points. We then reverse the color reconstruction of the colored marching cubes algorithm to derive colors for these key query point pairs, applying linear interpolation to assign colors to non-key query points spatially. The training approach aligns with the geometric coherent color representation strategy.
10 Color Ablation Study Details
We design a shape-agnostic strategy to display the superiority of our coherent 3D shape and color modeling. This strategy reverses the marching cubes algorithm to sample color values at uniformly distributed points within the query space, such as a $64\times 64\times 64$ resolution grid, as shown in Fig. 16. The difference between shape-agnostic strategy and geometric coherent color representation lies in the different methods of generating training data. We performed the marching cubes algorithm in reverse to identify the query coordinates $p_{1}$ and $p_{2}$ associated with each reconstructed object vertex. According to the definition of color reconstruction in the colored marching cubes algorithm:
@@ 10 Color Ablation Study Details
Figure 17: Unconditional 4D Generation Comparison. We highlight the parts for detailed comparison.
11 Visualization of Unconditional 4D Sequence Shape Generation
Fig. 17 provides a qualitative comparison of 4D sequences generated by HyperDiffusion and our method, illustrating that our model produces smoother surfaces with more detailed features, such as facial details. This highlights our model’s superior ability to generate 4D sequences with refined geometric details.

### v3

@@ 3.3 Spatio-Temporal Diffusion for Conditional 4D Volumetric Generation
$\mathcal{L}_{LDM}=\mathop{\mathbb{missing}}{E}_{\mathcal{M},f,\epsilon,t}\left[{\epsilon-\epsilon_{\theta}(\mathcal{M}^{t},t,F\{D\circ\tau_{i}(f_{i})\})}^{2}\right],$ (16)
where $D$ is a dropout operation enabling classifier-free guidance, $\tau_{i}(f_{i})$ is the $i$-th modality feature encoded by the task-specific encoder $\tau_{i}$, $\epsilon_{\theta}$ is the denoiser equipped with our HCSTA module, and $F$ refers to a simple concatenation. This approach not only ensures uniformity in the denoising process but also significantly reduces computational overhead.
@@ 4 Experiments
Data Preparation. For 3D geometry evaluation, we use car, chair, and airplane categories from ShapeNet [4]. For colored 3D objects evaluation, we use T-pose objects from 3DBiCar [23]. We use 16-frame animal sequences from DeformingThings4D (DT4D) [20] and 20-frame colored cartoon sequences from 3DBiCar (sample $30$k sequences via latent space interpolation) for both unconditional and conditional generation. DeformingThings4D [20] is a synthetic dataset containing $1,972$ animation sequences spanning 31 categories of humanoids and animals. However, the lack of color information and image annotations in DeformingThings4D makes it challenging to use for conditional 4D sequence generation. The 3DBiCar [23] dataset provides a script for sampling high-quality, topologically consistent 3D textured models of biped cartoon characters, manually crafted by professional artists, covering diverse identities, shapes, and textures across $15$ species and $4$ image styles. We also use $110$k subset and $323$k of Objaverse [7] for large scale conditional 4D generation. Objaverse is a large-scale 3D dataset containing both static and dynamic objects with more complex materials and geometric details, making it suitable for evaluating model performance on large-scale data. We select high-quality 4D sequences from the $110$k and $323$k subsets to train our method.
Both the 3DBiCar and Objaverse datasets provide image descriptions with masks relatively easily, but most of the data lack textual annotations. We employ Large Language Models (LLMs) for initial image descriptions, manually refined for accuracy in shape, color, and pose, with plans to release these captions post-acceptance. For conditioning, masked images were derived from provided masks and replicated across frames, while text conditions incorporated frame-order prefixes (e.g., ”$n$-th frame”) and were encoded with CLIP alongside image conditions. Training data included T-pose to target-pose transitions for model optimization.
Evaluation Metrics. We follow prior works [9, 5] in evaluating Minimum Matching Distance (MMD), Coverage (COV), and 1-Nearest-Neighbor Accuracy (1-NNA):
$\text{COV}(S_{g},S_{r})=\frac{|\{\arg\min\limits_{Y\in S_{r}}D(X,Y)|X\in S_{g}\}|}{|S_{r}|}$ (18)
[…]
where $D(X,K)$ is some kind of distance measure, and in the 1-NNA metric $N_{X}$ is a point cloud that is closest to $X$ in both generated and reference dataset:
$N_{X}=\arg\min\limits_{K\in S_{r}\cup S_{g}}D(X,K)$ (20)
where $S_{r}$ is the reconstructed 4D sequence, $S_{g}$ is the ground-truth 4D sequence. We also use a Chamfer Distance (CD) distance measure $D(X,Y)$ for computing these metrics in 3D, and report CD values multiplied by a constant $10^{2}$. To evaluate 4D shapes, we extend to the temporal dimension for $T$ frames:
$D(X,Y)=\frac{1}{T}\sum_{t=0}^{T-1}CD(X[t],Y[t]).$ (21)
For MMD and CD, lower is better; for COV, higher is better; for 1-NNA, 50% is the optimal. For 4D sequence evaluation, we average 3D metrics across frames. To assess inter-frame consistency, we compute CD-VAR and COV-VAR by calculating Chamfer Distance and COV variance across frames. For color evaluation, PSNR and SSIM are computed by sampling 16,384 surface points multiple times. We also use CLIP score (CLIP-S). All videos are rendered from Blender.
Implementation Details. The $\sigma$ is computed as $0.3\%$ of the original 4D sequence reconstruction loss. The loss weights $\lambda_{1}$ and $\lambda_{2}$ are set to $1$. For the shape auto-decoder $D_{\theta}$, we employ an 8-layer MLP to decode the SDF values, using a resolution of $64\times 64\times 64$ to construct a 6-layer 3D U-Net $U_{\phi}$ on DT4D and $128\times 128\times 128$ on 3DBiCar, ShapeNet, and Objaverse. For the color auto-decoder $C_{\theta}$, we use an 8-layer MLP. For denoising network, we use an 8-layer UNet augmented with HCSTA Layer.
@@ 4.1 Ablation Studies
Dimension of latents and loss weights. We provide an overview of the hyperparameter of our models in Tab. 4.1, including the size of the latent 4D representation $\mathcal{M}$ and loss weights. We evaluated the metrics and finally selecting the combination $\{\mathcal{M}\in\mathbb{R}^{N\times(256+256)},~{}\lambda_{1}=1.0,~{}\lambda_{2}=1.0\}$. Because balancing the L1 and SSIM losses aids the model in learning both the primary colors and the finer color details in a more balanced manner. Moreover, experimentation with dimensions larger than the current setting (256+256) was conducted. However, it is observed that when the scale of $\mathcal{M}$ exceeds 256 (as indicated in row 3 of Tab. 4.1), the data distribution becomes too sparse to support the diffusion generation for volumetric 4D sequence effectively.
@@
tableAblation of hyperparameter.
Category $U_{\phi}$ MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
$\times$ 0.012 0.007 69 2.791 36
3DBiCar $\checkmark$ 0.009 0.004 73 0.834 41
$\times$ 0.037 0.003 69 3.179 36
DT4D $\checkmark$ 0.030 0.001 76 2.91 43
[…]
tableAblation study of $U_{\phi}$.
@@
[…]
tableQuantitative comparison of Single Shape.
Finally, analysis of the remaining parameter settings (rows 4-5) reveals that an unbalanced weighting of loss terms significantly leads to performance degradation.
Refinement Module $U_{\phi}$. First, we compare 3D shape reconstruction quality among SDFusion [5], DeepSDF [26], and our method. SDFusion uses a 3D latent diffusion model with a VQVAE decoder for shape generation and a 2D diffusion model for texture rendering. DeepSDF decodes SDF values without color using an MLP decoder. We evaluate geometry accuracy on ShapeNet (car, chair, airplane) and 3DBiCar (T-pose) datasets. Tab. 4.1 shows our method outperforms SDFusion and DeepSDF in geometry generation and surpasses SDFusion in color generation, confirming that the refinement U-Net $U_{\phi}$ improves the quality of coarse SDF with a 3D U-Net, making our coherent 3D shape and color modeling learn more shape details. Then we ablate $U_{\phi}$ in 4D sequence generation, as demonstrated in Tab. 4.1, where $U_{\phi}$ improves all metrics because $U_{\phi}$ enhances the geometric performance of all frames within the 4D sequence, thereby indirectly promoting consistency across frames.
HCSTA module. Tab. 1 shows ablation studies on attention modulation, order, and unification. We first evaluate attention for shape-color (S), conditioned (C), and temporal (T) dimensions. Tab. 1 (rows 1-3) shows shape-color cross-attention contributes most. In rows 4-7, we permute the S, C, T orders, the S-C-T order performing best. Because our color field is geometry-aware, where accurate shape modeling facilitates precise subsequent generation of both color and motion. Finally, replacing the proposed hierarchical attention with a unified strategy (shared attention) leads to a significant drop (rows 8-9), our method outperforms it. Because our decoupled latent and attention design enables the model to explicitly learn accurate shape, color, and spatiotemporal relationships, leading to significant performance gains.
Type S C T MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
Module $\times$ $\checkmark$ $\checkmark$ 0.014 0.018 61 2.877 33
$\checkmark$ $\times$ $\checkmark$ 0.010 0.015 68 2.165 39
$\checkmark$ $\checkmark$ $\times$ 0.011 0.011 71 2.663 38
Order S-C-T 0.009 0.007 73 1.834 41
S-T-C 0.010 0.011 71 2.165 39
C-T-S 0.012 0.015 68 2.663 38
C-S-T 0.011 0.014 69 2.731 35
Uni-Stra. $\checkmark$ 0.013 0.018 67 3.002 34
$\times$ 0.009 0.007 73 1.834 41
Table 1: Ablation on attentions modulation, orders, and unifying.
Coherent 3D Shape and Color Modeling. We evaluate the impact of L1 loss, remeshing (R), structural similarity loss (SSIM), and a shape-agnostic approach (F) that assigns colors from a uniform distribution per object. This strategy reverses the marching cubes algorithm to sample color values at uniformly distributed points within the query space, such as a $64\times 64\times 64$ resolution grid:
@@
As evidenced in Tab. 5, integrating L1, R, and SSIM optimizes color fidelity: L1 ensures the overall color accuracy, SSIM improves local structural coherence, and R enhances vertex density for refined color detail. Our coherent 3D shape and color modeling constructs a continuous color field adaptive to varied geometries, ensuring shape-constrained color continuity and boosting accuracy. In contrast, shape-agnostic color learning severely weakens the correlation between color and geometry, leading to a significant drop in color fitting quality.
Figure 5: Color editing without shape alteration.
L1 R SSIM F PSNR SSIM
$\checkmark$ $\checkmark$ $\times$ $\times$ 29.15 0.79
$\checkmark$ $\times$ $\checkmark$ $\times$ 26.37 0.72
$\times$ $\checkmark$ $\checkmark$ $\times$ 28.76 0.75
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 9.64 0.64
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 34.51 0.85
[…]
tableQuantitative results of color ablation study.
Decoupled Color and Shape. We validate the efficacy of our decoupled shape and color framework by assessing fitting quality for combined and decoupled latents in Tab. 4.1. Conditional 4D generation quality and shape-color consistency are further evaluated through a user study in Tabs. 4.1, where participants compared our method to baselines across three metrics, selecting the superior approach per vote. Fig. 5 demonstrates that the decoupled representation allows color editing without altering shape. Results in Tabs. 4.1 consistently show that ours improves generation quality and enables independent color modifications. The explicit decoupling of shape and color allows for stable and independent optimization of structural and appearance components, thereby significantly enhancing intra-frame consistency by preventing color drifting and boosting inter-frame coherence by enabling smooth and geometry-aware temporal transitions.
[…]
4.2 Unconditional 4D Sequence Shape Generation
We first demonstrate the capability of our model to generate diverse objects and motions in unconditional 4D generation, then we showcase its ability to produce coherent sequences through latent matrixized 4D representation interpolation. We conduct evaluation on 3DBiCar and DeformingThings4D datasets, focusing on shape and temporal consistency. Tab. 2 presents quantitative comparisons of 4D sequence generation, demonstrating that our model consistently outperforms the voxel baseline and HyperDiffusion [9] (Hyper.) across all evaluated metrics. Because the voxel baseline and HyperDiffusion face challenges related to voxel resolution, which limits geometric precision, and high-dimensional sparsity in the weight space, which hindering HyperDiffusion from reconstructing high-quality 4D sequences, our method improves unconditional 4D generation by employing a matrixized 4D representation, which reduces the learning difficulty of denoiser during the diffusion process.
Fig. 7 provides a qualitative comparison of 4D sequences generated by HyperDiffusion and our method, illustrating that our model produces smoother surfaces with more detailed features, such as facial details. This highlights our superior ability to generate 4D sequences with refined geometric details. We also show the most similar objects in DT4D dataset to unconditional generation results in Fig.7, demonstrating the ability of our method to generate diverse 4D sequences.
Category Method MMD CD-VAR COV(%) COV-VAR 1-NNA(%)
Voxel 0.293 0.035 30 1.857 91
Hyper. [9] 0.044 0.011 52 2.138 69.3
3DBiCar Ours 0.009 0.007 73 1.834 41
Voxel 0.219 0.044 35 1.673 85
Hyper. [9] 0.155 0.013 45 1.735 62
DT4D Ours 0.030 0.001 76 2.91 43
Table 2: Quantitative comparison of 4D Sequence Generation.
Figure 6: Unconditional 4D Generation Comparison.